OpenStudy (anonymous):

find the value of h'(2) where h(x) = x^2/f(x), f(2)=3 and f ' (2)= -2 please show all calc involved

OpenStudy (accessdenied):

So, how would you find h'(x) ?

OpenStudy (anonymous):

would you plug in 2? no?

OpenStudy (anonymous):

write down the quotient rule, then make the replacements

OpenStudy (anonymous):

\[(\frac{g}{f})'=\frac{fg'-gf'}{f^2}\] with \[f(x)=f(x), f'(x)=f'(x), g(x)=x^2, g'(x)=2x\]

OpenStudy (anonymous):

is this clear? you should get \[h'(x)=\frac{2xf(x)-x^2f'(x)}{f^2(x)}\] then replace x by 2 to get your answer

OpenStudy (anonymous):

should i get 20/9

OpenStudy (anonymous):

i didn't do it. the 9 is right

OpenStudy (anonymous):

numerator is \[2\times 2\times 3+2^2\times 2\]

OpenStudy (anonymous):

yeah it is 20

OpenStudy (anonymous):

why isnt f^2(x) in the denominator 3^2 (2) which would be 18

OpenStudy (anonymous):

\[f(2)=3\] \[f^2(2)=3^2=9\]

OpenStudy (anonymous):

ok i got it. for a moment it looked like f^2 times (x)